# Speech-FT: Merging Pre-trained And Fine-Tuned Speech Representation Models For Cross-Task Generalization

### v2

@@ V-A Generality of Speech-FT across Models and Evaluation Tasks
We first apply Speech-FT to two widely used speech representation models that have distinct pre-training objectives [47] from HuBERT: wav2vec 2.0 [2], which is trained with a contrastive objective; and DeCoAR 2.0 [3], which is based on a generative approach.
The results are shown in Table VII. Speech-FT consistently improves performance across evaluation tasks for both models. In contrast, Stable-FT often results in degraded performance on tasks such as SID and ER. For example, when applied to DeCoAR 2.0, Speech-FT increases the SUPERB score from 736.77 (pre-trained model) to 781.77, whereas Stable-FT reduces it to 701.6. Similar results are observed for wav2vec 2.0, highlighting the generality of Speech-FT across models with different pre-training paradigms.
@@
We further evaluate Speech-FT on WavLM Base+ [5], a model that represents the state of the art among base-sized (Transformers of 12 layers) speech representation models. Notably, WavLM Base+ is pre-trained on 94,000 hours of unlabelled speech, in contrast to the 960 hours used for HuBERT, wav2vec 2.0 and DeCoAR 2.0, resulting in significantly more general speech representations. Although the pre-trained model already demonstrates strong performance, Speech-FT is still able to deliver further improvements. In particular, the PER is reduced from 3.93% to 3.42% on PR, corresponding to a relative improvement of 12.97%. Although there is a minor degradation on SID, the overall SUPERB score still improves from 942.96 to 945.5. These results indicate that Speech-FT remains effective even on models that have been pre-trained on significantly larger datasets.
Speech-FT remains effective on full SUPERB evaluation.
We further evaluate Speech-FT on all ten tasks in the SUPERB benchmark, reporting the comprehensive SUPERBFULL score. Unlike the SUPERBS scores used elsewhere in the paper, SUPERBFULL encompasses the complete set of tasks of SUPERB when computing Equation 9. In addition to PR, SID, ER, and SF, the full set includes: Automatic Speech Recognition (ASR), which transcribes audio into text and is evaluated using word error rate (WER); Keyword Spotting (KS), which detects target keywords and is evaluated by classification accuracy; Query by Example (QbE), which retrieves matching audio segments using maximum term-weighted value (MTWV); Automatic Speaker Verification (ASV), which verifies speaker identity using equal error rate (EER); Speaker Diarization (SD), which segments who spoke when using diarization error rate (DER); and Intent Classification (IC), which predicts utterance intent based on classification accuracy.
@@
We select HuBERT and WavLM Base+ as representative models. The results are presented in Table VIII. Speech-FT significantly enhances the performance of HuBERT, raising its SUPERBFULL score from 815.47 to 874.51. Consistent improvements are observed across content-related tasks (e.g., ASR: 6.38% WER to 5.75%), speaker-related tasks (e.g., ASV: 5.56% EER to 5.27%), and semantic-related tasks (e.g., IC: 98.02% to 98.97%).
@@
For WavLM Base+, whose pre-trained model already exhibits strong performance, Speech-FT leads to minor drops on speaker-related tasks such as ASV and SD, but still yields gains on most tasks. Notably, the overall SUPERBFULL score increases from 946.95 to 950.60. These findings confirm that Speech-FT remains effective under extensive evaluation settings.
@@
TABLE X: Comparison of Speech-FT with alternative fine-tuning strategies. Experiments are conducted on phoneme classification (PC) using the TIMIT dataset and speaker identification (SID) using VoxCeleb1. “Weight-Space Reg.” denotes fine-tuning with weight-space regularization. “LoRA” indicates LoRA fine-tuning. “Early Checkpoint (20%)” denotes the model snapshot saved at 20% of the total fine-tuning steps. The background color highlights cases where the evaluation task is included in the fine-tuning tasks.
@@ V-B Stable-FT Enhances Cross-Task Generalization in Speech-FT
To understand the impact of Stable-FT in Speech-FT, we conduct an ablation study by removing it from the Speech-FT framework. This analysis is conducted on phoneme classification (PC) on TIMIT and speaker identification (SID) on VoxCeleb1. As shown in Table IX, removing Stable-FT significantly impairs the cross-task generalization ability of the model.
@@ V-C Comparison to Baselines That Constrain Weight Deviation During Fine-tuning
Speech-FT outperforms weight-space regularization. Weight-space regularization offers a straightforward alternative to our interpolation-based approach by explicitly constraining parameter updates. During fine-tuning, we add a regularization term $\lambda\|\theta^{\prime}-\theta_{0}\|^{2}_{F}$ to penalize deviation from the pre-trained weights. The hyperparameter $\lambda$ is selected from $\{10^{-2},10^{-4},10^{-6}\}$, and we report the best-performing configuration.
As shown in Table X, weight-space regularization improves performance on tasks aligned with the fine-tuning objective, but consistently degrades performance on those that are not. For instance, when fine-tuned on PC, models trained with regularization perform worse than both Speech-FT and the pre-trained model on unrelated tasks such as SID and ER. SID accuracy drops to 80.67% and ER accuracy to 64.84%, both lower than the pre-trained scores of 81.86% and 64.99%, respectively. Similarly, when fine-tuned on SID, PER increases to 9.94% on PR, and ER accuracy drops to 63.88%, again underperforming both the pre-trained model and Speech-FT.
@@
TABLE XI: The impact of hyperparameter $\alpha$ when fine-tuning with phoneme classification on TIMIT.
$\alpha$ PR SID ER SF $\text{SUPERB}_{\text{S}}\uparrow$
PER%$\downarrow$ ACC%$\uparrow$ ACC%$\uparrow$ F1$\uparrow$ CER%$\downarrow$
0.0 5.17 81.86 64.99 88.54 24.70 870.20
0.10 4.99 81.19 65.33 88.69 24.01 874.68
0.25 4.76 81.78 65.48 88.65 24.05 877.66
0.50 5.15 80.24 64.28 87.03 27.22 843.75
0.75 6.63 75.34 63.10 86.55 28.47 810.79
1.0 10.34 66.34 61.25 83.5 33.82 726.64
Speech-FT outperforms LoRA fine-tuning. To further contextualize the effectiveness of Speech-FT, we compare it with LoRA fine-tuning [17] as an additional baseline. LoRA was originally proposed to improve fine-tuning efficiency by introducing low-rank adaptation modules while keeping the pre-trained weights fixed. It also serves as a relevant baseline for Speech-FT, as it similarly aims to minimize changes to the pre-trained model. However, while LoRA constrains changes in weight space, Speech-FT instead constrains changes in feature space. We follow the implementation of prior work [48] for LoRA fine-tuning on speech representation models.
As shown in Table X, LoRA consistently underperforms Speech-FT across all tasks, with particularly large performance drops on unrelated evaluation tasks. For instance, when fine-tuned on PC, LoRA achieves 71.43% accuracy on SID, 63.34% on ER, and a SUPERB score of 824.72. In comparison, Speech-FT achieves 81.78%, 65.48%, and 877.66, respectively.
@@
Speech-FT outperforms early checkpoints during fine-tuning. We also examine whether stopping the fine-tuning process early could help preserve cross-task generalization. Specifically, we compare Speech-FT with checkpoints obtained at 20%, 40%, and 60% of the total fine-tuning steps. These checkpoints represent intermediate stages of fine-tuning, where the model has not yet fully adapted its parameters and thus remains relatively close to the pre-trained model in weight space.
Among the 20%, 40%, and 60% checkpoints, the 20% checkpoint yields the most favorable results, which we report in Table X. Although early checkpoints could introduce fewer deviations in weight space compared to full fine-tuning, they still lead to degraded performance across all evaluation tasks. When fine-tuned on PC, the 20% checkpoint yields a SUPERB score of 790.44, while SID accuracy drops from 81.86% (pre-trained model) to 71.80%. Similarly, when fine-tuned on SID, the SUPERB score decreases to 736.06, accompanied by an increase in PER on PR from 5.17% (pre-trained model) to 19.75%. These findings suggest that early stopping alone is insufficient to preserve cross-task generalization.
Figure 2: Feature similarity with the pre-trained model. (Left) Effect of $\alpha$ on the cosine similarity between Speech-FT and pre-trained features. (Middle) Cosine similarity between the pre-trained features and those from Speech-FT, weight-space regularization (“Weight-Space Reg.”), LoRA fine-tuning (“LoRA”), early checkpoint during fine-tuning (“Early Checkpoint”), and feature-space regularization (“Feature-Space Reg.”). (Right) L2 distance between the pre-trained features and those from Speech-FT, weight-space regularization, LoRA fine-tuning, early checkpoint during fine-tuning, and feature-space regularization.
@@ V-D Speech-FT Preserves Feature Similarity Better Than Alternatives
In the previous section, we showed that Speech-FT achieves better cross-task generalization than other alternatives that aim to minimize deviations in weight space. To better understand why Speech-FT leads to improved cross-task generalization, we analyze the feature similarity between Speech-FT and the pre-trained model in this section. Motivated by the findings, we further introduce a feature-space regularization baseline to investigate whether constraining representational drift via a regularization term during fine-tuning can help preserve cross-task generalization. All analyses are conducted on the TIMIT dataset for phoneme classification (PC).
Weight-space interpolation effectively restores feature similarity to the pre-trained model. Table XI shows how the performance of speech representations evolves as the interpolation hyperparameter $\alpha$ decreases from 1.0 (fully fine-tuned model) to 0.0 (pre-trained model). Fine-tuning without weight-space interpolation ($\alpha=1.0$) compromises the cross-task generalization ability of the speech representation. However, as $\alpha$ decreases, the SUPERB score gradually improves, reaching its peak at $\alpha=0.25$. As shown in the left panel of Figure 2, smaller values of $\alpha$ also lead to higher feature similarity with the pre-trained model. This helps explain the observed improvement in cross-task generalization when using lower values of $\alpha$.
TABLE XII: Comparison of Speech-FT and feature-space regularization (“Feature-Space Reg.”). We use phoneme classification on TIMIT as the fine-tuning task for this experiment.
@@
Speech-FT exhibits higher feature similarity to the pre-trained model than other alternatives
We further compare how similar the features of Speech-FT and other alternatives are to those of the pre-trained model. Although Speech-FT results in a larger change in weight space compared to weight-space regularization—as shown in Figure 3—it achieves much higher similarity at the feature level as shown in the middle panel of Figure 2. The right panel of Figure 2 further supports this observation using L2 distance as an additional measure, where Speech-FT again produces features more similar to those of the pre-trained model. It is worth noting that the regularization strength for weight-space regularization is a tunable hyperparameter, and all analyses in Figures 2 and 3 are based on its best-performing configuration.
These findings highlight a key difference between Speech-FT and weight-space regularization. While weight-space regularization could keep the fine-tuned model close to the pre-trained model in parameter space, it fails to preserve feature similarity. In contrast, Speech-FT, despite allowing more parameter changes, is better at maintaining the original feature space. This helps explain the results in Table X, where weight-space regularization degrades cross-task generalization more severely than Speech-FT. The reduced feature similarity in regularized models likely contributes to their weaker performance on tasks unrelated to the fine-tuning objective.
@@
Speech-FT outperforms feature-space regularization.
@@
As shown in the middle and right panels of Figure 2, feature-space regularization successfully maintains a high degree of similarity to the pre-trained model across all layers, as measured by both cosine similarity and L2 distance. This suggests that the feature-space regularization objective is effective in constraining the representations to remain close to those learned during pre-training.
Correspondingly, Table XII shows that this method could help preserve cross-task generalization ability. Specifically, feature-space regularization achieves 81.72% accuracy on SID and 65.01% accuracy on ER, which are comparable to the performance of the pre-trained model. Compared to the results of weight-space regularization reported in Table X, which achieve 80.67% on SID and 64.84% on ER, feature-space regularization yields notable improvements on tasks unrelated to the fine-tuning objective.

### v3

@@ V-A1 Speech-FT is effective across different speech representation models
We first apply Speech-FT to two widely used speech representation models that have distinct pre-training objectives [mohamed2022self] from HuBERT: wav2vec 2.0 [baevski2020wav2vec], which is trained with a contrastive objective; and DeCoAR 2.0 [ling2020decoar], which is based on a generative approach. The results are shown in Table V. Speech-FT consistently improves SUPERB score for both models. For instance, on DeCoAR 2.0, it increases the score from 736.77 (pre-trained) to 781.77, while Stable-FT decreases it to 701.60. Similarly, on wav2vec 2.0, Speech-FT raises the score from 837.08 to 864.79, whereas Stable-FT lowers it to 824.82.
@@
We further evaluate Speech-FT on WavLM Base+ [chen2022wavlm], a model that represents the state of the art among base-sized (Transformers of 12 layers) speech representation models. Notably, WavLM Base+ is pre-trained on 94,000 hours of unlabelled speech, in contrast to the 960 hours used for HuBERT, wav2vec 2.0 and DeCoAR 2.0, resulting in significantly more general speech representations. Although the pre-trained model already demonstrates strong performance, Speech-FT is still able to deliver further improvements. In particular, the overall SUPERB score improves from 942.96 to 945.50.
V-A2 Speech-FT remains effective on full SUPERB evaluation
We further evaluate Speech-FT on all ten tasks in the SUPERB benchmark, reporting the comprehensive SUPERBFULL score. Unlike the SUPERBS scores used elsewhere in the paper, SUPERBFULL encompasses the complete set of tasks of SUPERB when computing Equation 5. The task definitions and evaluation metrics follow the original SUPERB paper [yang2021superb].
@@ V-A2 Speech-FT remains effective on full SUPERB evaluation
We select HuBERT and WavLM Base+ as representative models. The results are presented in Table VI. Speech-FT significantly enhances the performance of HuBERT, raising its SUPERBFULL score from 815.47 to 874.51. Consistent improvements are observed across content-related tasks (e.g., ASR: 6.38% WER to 5.75%), speaker-related tasks (e.g., ASV: 5.56% EER to 5.27%), and semantic-related tasks (e.g., IC: 98.02% to 98.97%).
@@
For WavLM Base+, whose pre-trained model already exhibits strong performance, Speech-FT leads to minor drops on speaker-related tasks such as ASV and SD, but still yields gains on most tasks. For example, the PER is reduced from 3.94% to 3.42% on PR, corresponding to a relative improvement of 13.2%. Additionally, the overall SUPERBFULL score increases from 946.95 to 950.60. These findings confirm that Speech-FT remains effective under extensive evaluation settings.
@@
TABLE VIII: Comparison of Speech-FT with alternative fine-tuning strategies. Experiments are conducted on phoneme classification (PC) using the TIMIT dataset and speaker identification (SID) using VoxCeleb1. “Weight-Space Reg.” denotes fine-tuning with weight-space regularization. “LoRA” indicates LoRA fine-tuning [hu2022lora]. “DoRA” indicates DoRA fine-tuning [liu2024dora]. “Early Checkpoint (20%)” denotes the model snapshot saved at 20% of the total fine-tuning steps. The background color highlights cases where the evaluation task is included in the fine-tuning tasks.
@@ V-B Stable-FT Enhances Cross-Task Generalization in Speech-FT
To understand the impact of Stable-FT in Speech-FT, we conduct an ablation study by removing it from the Speech-FT framework. This analysis is conducted on phoneme classification (PC) on TIMIT and speaker identification (SID) on VoxCeleb1. As shown in Table VII, removing Stable-FT significantly impairs the cross-task generalization ability of the model.
@@ V-C Comparison to Baselines That Constrain Weight Deviation During Fine-tuning
V-C1 Speech-FT outperforms weight-space regularization
Weight-space regularization offers a straightforward alternative to our interpolation-based approach by explicitly constraining parameter updates. During fine-tuning, we add a regularization term $\lambda\|\theta^{\prime}-\theta_{0}\|^{2}_{F}$ to penalize deviation from the pre-trained weights. The hyperparameter $\lambda$ is selected from $\{10^{-2},10^{-4},10^{-6}\}$, and we report the best-performing configuration.
As shown in Table VIII, weight-space regularization improves performance on tasks aligned with the fine-tuning objective, but consistently degrades performance on those that are not. For instance, when fine-tuned on PC, models trained with regularization perform worse than both Speech-FT and the pre-trained model on unrelated tasks such as SID and ER. SID accuracy drops to 80.67% and ER accuracy to 64.84%, both lower than the pre-trained scores of 81.86% and 64.99%, respectively. Similarly, when fine-tuned on SID, PER increases to 9.94% on PR, and ER accuracy drops to 63.88%, again underperforming both the pre-trained model and Speech-FT.
@@ V-C1 Speech-FT outperforms weight-space regularization
TABLE IX: The impact of hyperparameter $\alpha$ when fine-tuning on phoneme classification (TIMIT), speaker identification (VoxCeleb1), and emotion recognition (IEMOCAP).
$\alpha$ PC on TIMIT SID on Voxceleb1 ER on IEMOCAP
[…]
0.0 870.20 870.20 870.2
0.10 874.68 876.89 932.59
0.25 877.66 881.07 1010.29
0.50 843.75 810.36 1007.38
0.75 810.79 721.54 983.72
1.0 726.64 651.42 908.86
V-C2 Speech-FT outperforms parameter-efficient fine-tuning methods
To further contextualize the effectiveness of Speech-FT, we compare it with LoRA fine-tuning [hu2022lora] as an additional baseline. LoRA was originally proposed to improve fine-tuning efficiency by introducing low-rank adaptation modules while keeping the pre-trained weights fixed. It also serves as a relevant baseline for Speech-FT, as it similarly aims to minimize changes to the pre-trained model. However, while LoRA constrains changes in weight space, Speech-FT instead constrains changes in feature space. We follow the implementation of prior work [chen2023explore] for LoRA fine-tuning on speech representation models.
As shown in Table VIII, LoRA consistently underperforms Speech-FT across all tasks, with particularly large performance drops on unrelated evaluation tasks. For instance, when fine-tuned on PC, LoRA achieves 71.43% accuracy on SID, 63.34% on ER, and a SUPERB score of 824.72. In comparison, Speech-FT achieves 81.78%, 65.48%, and 877.66, respectively.
@@ V-C2 Speech-FT outperforms parameter-efficient fine-tuning methods
We also experiment with DoRA fine-tuning [liu2024dora], a recently proposed PEFT method that improves on LoRA by decomposing weights into magnitude and direction. While DoRA achieves slightly higher SUPERB scores than LoRA, it nevertheless suffers from the same drawback and shows severely degraded cross-task generalization compared to Speech-FT.
V-C3 Speech-FT outperforms early checkpoints during fine-tuning
We also examine whether stopping the fine-tuning process early could help preserve cross-task generalization. Specifically, we compare Speech-FT with checkpoints obtained at 20%, 40%, and 60% of the total fine-tuning steps. These checkpoints represent intermediate stages of fine-tuning, where the model has not yet fully adapted its parameters and thus remains relatively close to the pre-trained model in weight space.
Among the 20%, 40%, and 60% checkpoints, the 20% checkpoint yields the most favorable results, which we report in Table VIII. Although early checkpoints could introduce fewer deviations in weight space compared to full fine-tuning, they still lead to degraded performance across all evaluation tasks. When fine-tuned on PC, the 20% checkpoint yields a SUPERB score of 790.44, while SID accuracy drops from 81.86% (pre-trained model) to 71.80%. Similarly, when fine-tuned on SID, the SUPERB score decreases to 736.06, accompanied by an increase in PER on PR from 5.17% (pre-trained model) to 19.75%. These findings suggest that early stopping alone is insufficient to preserve cross-task generalization.
Figure 2: Feature similarity with the pre-trained model. (Top) Effect of $\alpha$ on the cosine similarity between Speech-FT and pre-trained features. (Bottom) Cosine similarity between the pre-trained features and those from Speech-FT, weight-space regularization (“Weight-Space Reg.”), LoRA fine-tuning (“LoRA”), early checkpoint during fine-tuning (“Early Checkpoint”), and feature-space regularization (“Feature-Space Reg.”).
@@ V-D Speech-FT Preserves Feature Similarity Better Than Alternatives
In the previous section, we showed that Speech-FT achieves better cross-task generalization than other alternatives that aim to minimize deviations in weight space. To better understand why Speech-FT leads to improved cross-task generalization, we analyze the feature similarity between Speech-FT and the pre-trained model in this section. Motivated by the findings, we further introduce a feature-space regularization baseline to investigate whether constraining representational drift via a regularization term during fine-tuning can help preserve cross-task generalization. Finally, we provide a brief theoretical explanation of Speech-FT. All analyses are conducted on the TIMIT dataset for phoneme classification (PC).
V-D1 Weight-space interpolation effectively restores feature similarity to the pre-trained model
Table IX shows how the performance of speech representations evolves as the interpolation hyperparameter $\alpha$ decreases from 1.0 (fully fine-tuned model) to 0.0 (pre-trained model). Fine-tuning without weight-space interpolation ($\alpha=1.0$) compromises the cross-task generalization ability of the speech representation. However, as $\alpha$ decreases, the SUPERB score gradually improves, reaching its peak at $\alpha=0.25$. As shown in the top panel of Figure 2, smaller values of $\alpha$ also lead to higher feature similarity with the pre-trained model. This helps explain the observed improvement in cross-task generalization when using lower values of $\alpha$.
TABLE X: Comparison of Speech-FT and feature-space regularization (“Feature-Space Reg.”). We use phoneme classification on TIMIT as the fine-tuning task for this experiment.
@@ V-D1 Weight-space interpolation effectively restores feature similarity to the pre-trained model
V-D2 Speech-FT exhibits higher feature similarity to the pre-trained model than other alternatives
We further compare how similar the features of Speech-FT and other alternatives are to those of the pre-trained model. Although Speech-FT results in a larger change in weight space compared to weight-space regularization—as shown in Figure 3—it achieves much higher similarity at the feature level [^6] In addition to cosine similarity, we also measured the L2 distance between features, and the overall trend remains consistent. as shown in the bottom panel of Figure 2. It is worth noting that the regularization strength for weight-space regularization is a tunable hyperparameter, and all analyses in Figures 2 and 3 are based on its best-performing configuration.
These findings highlight a key difference between Speech-FT and weight-space regularization. While weight-space regularization could keep the fine-tuned model close to the pre-trained model in parameter space, it fails to preserve feature similarity. In contrast, Speech-FT, despite allowing more parameter changes, is better at maintaining the original feature space. This helps explain the results in Table VIII, where weight-space regularization degrades cross-task generalization more severely than Speech-FT. The reduced feature similarity in regularized models likely contributes to their weaker performance on tasks unrelated to the fine-tuning objective. Moreover, our layer-wise analysis in Figure 2 shows that Speech-FT consistently preserves higher similarity in the middle and upper layers, which are known to encode richer phonetic information [pasad2021layer]. This pattern suggests that Speech-FT may better retain phonetic knowledge from pre-training, consistent with its strong performance on the PR task in Table VIII.
@@ V-D2 Speech-FT exhibits higher feature similarity to the pre-trained model than other alternatives
V-D3 Speech-FT outperforms feature-space regularization
@@
As shown in the bottom panel of Figure 2, feature-space regularization successfully maintains a high degree of similarity to the pre-trained model across all layers, as measured by both cosine similarity and L2 distance. This suggests that the feature-space regularization objective is effective in constraining the representations to remain close to those learned during pre-training.
Correspondingly, Table X shows that this method could help preserve cross-task generalization ability. Specifically, feature-space regularization achieves 81.72% accuracy on SID and 65.01% accuracy on ER, which are comparable to the performance of the pre-trained model. Compared to the results of weight-space regularization reported in Table VIII, which achieve 80.67% on SID and 64.84% on ER, feature-space regularization yields notable improvements on tasks unrelated to the fine-tuning objective.
@@
V-D4 Theoretical explanation of Speech-FT
TABLE XI: LLFC verification under Speech-FT. Endpoints ($\alpha\in\{0,1\}$) omitted for brevity. Features are orthogonally aligned to the pre-trained endpoint. Higher $R^{2}$ indicates stronger layerwise linearity, and regression coefficients $(b_{0},b_{1})$ closer to the theoretical $(1-\alpha,\alpha)$ indicate better agreement with the predicted linear mixture.
Method $\boldsymbol{\alpha}$ Linear mixture fit
$\mathbf{b_{0}}$ $\mathbf{b_{1}}$ $\mathbf{R^{2}}\uparrow$
Speech-FT 0.25 0.782 0.220 0.985
0.50 0.491 0.507 0.980
0.75 0.219 0.779 0.989
Speech-FT - Stable fine-tuning 0.25 0.831 0.185 0.985
0.50 0.536 0.474 0.977
0.75 0.239 0.763 0.986
The theoretical basis of Speech-FT can be explained through the lens of Layerwise Linear Feature Connectivity (LLFC) [zhou2023going]. LLFC was originally defined for image classification models fine-tuned from the same pre-trained initialization, where internal representations evolve almost linearly along the interpolation path. In our setting, one endpoint is the pre-trained model $\theta_{0}$ and the other is its fine-tuned variant $\theta^{\prime}$. Since $\theta^{\prime}$ is directly obtained from $\theta_{0}$, the same LLFC principle applies: the per-layer features along the interpolation path $F_{\ell}(\hat{\theta}(\alpha))\approx(1-\alpha)F_{\ell}(\theta_{0})+\alpha F_{\ell}(\theta^{\prime})$ are well approximated by the linear mixture of the two endpoints’ features, indicating that representation geometry is largely preserved.
To verify this in practice, we evaluate LLFC on a development set by extracting per-layer features $F_{\ell}(\cdot)$ from $\theta_{0}$, $\theta^{\prime}$, and interpolated checkpoints $\hat{\theta}(\alpha)$ ($\alpha\in\{0.25,0.50,0.75\}$). After aligning these features to $F_{\ell}(\theta_{0})$ via Orthogonal Procrustes [schonemann1966procrustes], thereby removing arbitrary rotations between representation spaces, we compute a ridge regression [hoerl1970ridge] of $F_{\ell}(\hat{\theta}(\alpha))$ on $\{F_{\ell}(\theta_{0}),F_{\ell}(\theta^{\prime})\}$, yielding coefficients $(b_{0},b_{1})$ and the coefficient of determination $R^{2}$, with higher values indicating a better fit. As summarized in Table XI, Speech-FT exhibits high $R^{2}$ values and regression coefficients close to the theoretical $(1-\alpha,\alpha)$, confirming that it satisfies LLFC. Additionally, the stable fine-tuning step appears to play a crucial role in supporting this behavior.
This means interpolation explicitly pulls the features back toward pre-training directions, thereby restoring similarity to the original representations. In contrast, weight-space regularization only constrains parameter distance without ensuring that the resulting models lie in a representation-preserving region. As a result, it does not guarantee recovery of the pre-training feature geometry.
While LLFC explains the mechanism for restoring feature similarity, the framework of Task Arithmetic [ilharco2022editing] provides another perspective for why task-specific knowledge is simultaneously retained. This framework treats the changes from fine-tuning as a task vector—the arithmetic difference between the fine-tuned and pre-trained model weights—which encapsulates the knowledge learned for a specific task. By defining this task vector as $\tau=\theta^{\prime}-\theta_{0}$, our interpolation is equivalent to: $\hat{\theta}(\alpha)=\theta_{0}+\alpha\tau$. This clarifies that Speech-FT operates by adding a scaled version of the task-specific knowledge ($\tau$) to the pre-trained base model ($\theta_{0}$). This process inherently preserves the general capabilities of $\theta_{0}$ while controllably injecting the specialized knowledge contained within $\tau$.